OpenStudy (harley1312):

Can someone plz help me? If X = 6 cm and Y = 8 cm, what is the length of Z?

OpenStudy (anonymous):

is there a pic ?

OpenStudy (harley1312):

Here is one

OpenStudy (anonymous):

Y\[Z ^{2} = X^2 + Y^2\] \[Z^2 = 36 + 64\] \[z^2 = 100\] \[z = \sqrt{100} = 10\]

OpenStudy (anonymous):

answer : B

OpenStudy (harley1312):

Thank you so much i gave u medal :)